Given that sinx = 3/5 and cosy = 7/25, with x and y both being Quadrant I angles, find tan(x/2)
Answer by ikleyn(53938)

.
Given that sinx = 3/5 and cosy = 7/25, with x and y both being Quadrant I angles, find tan(x/2)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The condition and the question are not balanced.